Shirley and Jim, working together, can clean the house in 8 hours. Working alone, Jim takes twice as long as Shirley. How long does it take Shirley to clean the house alone?
Found 2 solutions by greenestamps, ikleyn:
Answer by greenestamps(13209)



If Jim takes twice as long as Shirley to clean the house, then when working together Jim does 1/3 of the work and Shirley does 2/3.

In taking 8 hours to do the job together, Jim does 1/3 of the work; that means it would take him 8*3=24 hours to do the job alone.

And since he alone takes twice as long a Shirley, Shirley would take 12 hours to do the job alone.


Answer by ikleyn(52878)

.

It is the same as if three instances of Jim complete the job in 8 hours.


Hence, one single Jim can do the job in 3*8 = 24 hours.


24 hours is twice as long as 12 hours - hence, the time for Shirley is 24/2 = 12 hours to complete the job alone.   ANSWER